# From Reasoning to Learning: A Survey on Hypothesis Discovery and Rule Learning with Large Language Models

### v1

@@ 1 Introduction
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=6.5em,font=,, where level=2text width=10.5em,font=,, where level=3text width=13.5em,font=,, where level=4text width=10.5em,font=,, where level=5text width=6.5em,font=,, [ Reasoning for Hypothesis Discovery, ver [ Hypothesis
Generation w/
Abduction § 4 [ Method § 4.1 [ Natural Language Hypothesis [ Prompt Method [ Wiegreffe et al. (2022); Qi et al. (2024); Yang et al. (2024a), leaf, text width=32em ] ] [ RAG Method [ Hu et al. (2024); Yang et al. (2025; 2024b); Xiong et al. (2024)
Chai et al. (2024), leaf, text width=32em ] ] [ Human supported [ Zhao et al. (2024); Si et al. (2024); Pu et al. (2025), leaf, text width=32em ] ] ] [ Formal Language Hypothesis [ W/ Formal observation [ Young et al. (2022); Nguyen et al. (2023), leaf, text width=32em ] ] [ W/ NL observation [ Cheng et al. (2024); Wang et al. (2024); Qiu et al. (2024)
il Lee et al. (2025); Movva et al. (2025), leaf, text width=32em ] ] ] ] [ Evaluation § 4.2 [ Natural Language Hypothesis [ Prediction Based [ Sinha et al. (2019); Weston et al. (2015); Balepur et al. (2024)
Shi et al. (2023); Wang et al. (2025), leaf, text width=32em ] ] [ Ground Truth Based [ Yang et al. (2024a); Qi et al. (2024); Movva et al. (2025)
Yang et al. (2024b), leaf, text width=32em ] ] [ Human Evaluation [ Zhao et al. (2024); Yang et al. (2024b); Hu et al. (2024); Qi et al. (2024)
Yang et al. (2025), leaf, text width=32em ] ] ] [ Formal Language Hypothesis [ Prediction Based [ Rule (2020); Chollet (2019); Liu et al. (2024); Li et al. (2025a)
Chen et al. (2024), leaf, text width=32em ] ] [ Ground Truth Based [ Bowen et al. (2024); Hua et al. (2025); Young et al. (2022), leaf, text width=32em ] ] ] ] ] [ Hypothesis
Application w/
Deduction § 5 [ Method § 5.1 [ Natural Language Hypothesis [ LLM as FL Parser [ Pan et al. (2023); Olausson et al. (2023); Kalyanpur et al. (2024), leaf, text width=32em ] ] [ Fine-Tune Based [ Sun et al. (2024), leaf, text width=32em ] ] [ Prompt-Based [ Sun et al. (2024); Cai et al. (2025); Ling et al. (2023), leaf, text width=32em ] ] ] [ Formal Language Hypothesis [ N/A ] ] ] [ Evaluation § 5.2 [ Natural Language Hypothesis [ Prediction Based [ Yu et al. (2024b); Mu et al. (2024); Cai et al. (2025); Sun et al. (2024), leaf, text width=32em ] ] ] [ Formal Language Hypothesis [ N/A ] ] ] ] [ Hypothesis
Validation w/
Induction § 6 [ Method § 6.1 [ Natural Language Hypothesis [ Prompt-Based [ Lampinen et al. (2022); Sun et al. (2024), leaf, text width=32em ] ] [ Fine-Tune Based [ Zhao et al. (2023); Chan et al. (2023); Sun et al. (2024), leaf, text width=32em ] ] ] [ Formal Language Hypothesis [ Prompt-Based [ He & Lu (2024), leaf, text width=32em ] ] ] ] [ Evaluation § 6.2 [ Natural Language Hypothesis [ True/False Based [ Lampinen et al. (2022); Rudinger et al. (2020); Zhang et al. (2025), leaf, text width=32em ] ] [ Selection Based [ Bhagavatula et al. (2020); Jiang et al. (2023); Del & Fishel (2023), leaf, text width=32em ] ] ] [ Formal Language Hypothesis [ Ground Truth Based [ He & Lu (2024), leaf, text width=32em ] ] ] ] ] [ Entire
Hypothesis
Discovery § 7 [ Passive Hypothesis Discovery § 7.1 , text width=37.8em [ Zhu et al. (2024); Zhou et al. (2024), leaf, text width=32em ] ] [ Proactive Hypothesis Discovery § 7.2, text width=37.8em [ Li et al. (2024); Jung et al. (2022); Seals & Shalin (2024)
Conti et al. (2024), leaf, text width=32em ] ] [ Hypothesis Discovery with Real World Simulation § 7.3, text width=37.8em [ Xu et al. (2023); Wang et al. (2022a); Jansen et al. (2024)
He et al. (2024), leaf, text width=32em ] ] ] ]
@@ 2.2 Reasoning
Deductive Reasoning is the process of logically deriving specific conclusions from general hypotheses or rules. If the initial hypotheses are true, deduction guarantees that the derived conclusions must also be true. For instance, from the general rule “All swans are white” and the observation “This bird is a swan,” we logically conclude “This bird must be white”. While traditional deductive reasoning tasks, such as instruction-following and standard problem-solving, have been extensively studied with LLMs (Pan et al., 2023; Wei et al., 2022; Liu et al., 2025; Huang & Chang, 2023), deductive reasoning in the context of hypothesis discovery poses unique challenges. Specifically, it emphasizes inferential rule-following, requiring models to consistently apply hypotheses or rules to derive new and potentially unfamiliar conclusions, even when these hypotheses are counterfactual, unfamiliar, or incorrect. For example, when a flawed hypothesis is introduced in an unfamiliar domain, inferential rule-following requires us to strictly derive its predicted consequence, even if that consequence itself is incorrect. By comparing this consequence with experimental data, we can directly assess the hypothesis’s validity and guide its revision. Conversely, if the deductive process is unreliable, we may overlook real contradictions and thus retain invalid hypotheses or discard valid ones. Indeed, recent work shows that although LLMs can demonstrate strong deductive performance on in-distribution tasks, they rely heavily on surface-level pattern matching and fail to generalize their inferential rule-following to novel or counterfactual scenarios (Pu et al., 2025; Mirzadeh et al., 2024; Kang et al., 2024; Yan et al., 2025).
@@ 2.3 Rule Representation: Formal Language vs Natural Language
Natural Language English: ‘Sam is a dragon” English commonsense of ‘dragon” Sam is dangerous
Chinese: ‘Sam is a dragon” Chinese commonsense of ‘dragon” Sam brings good fortune and a bountiful harvest
Formal Language FOL: $\mathrm{Dragon}(\mathrm{Sam})$ $\forall x(\mathrm{Dragon}(x)\to\mathrm{Fly}(x))$ $\mathrm{Fly(Sam)}$
Code: Sam = Dragon() Class Dragon:      def fly(self):       … Sam.fly()
There are many ways to represent hypotheses and rules, which we broadly divide into two categories: formal languages (FL) and natural languages (NL). Formal languages, such as first-order logic and programming languages, are systematic and rule-bound. After real-world entities are encoded as explicit literals, precise inference rules yield provably correct and sound conclusions, making these systems well suited to deductive reasoning. Yet the encoding process strips away many subtle semantic relationships and commonsense knowledge, limiting the system’s ability to handle the creative, defeasible reasoning required for abduction and induction (McCarthy & Hayes, 1981; Reiter, 1980; Hanks & McDermott, 1987; Liu et al., 2025; Yu et al., 2024a; Huang & Chang, 2023). Natural language preserves those nuances and aligns more closely with human cognition, so it is better suited to abductive and inductive tasks. However, its meanings are implicit and context-dependent, making it difficult to define a deterministic reasoning pipeline and reducing the reliability of the resulting inferences (See in Table  1). Accordingly, the following sections treat formal-language and natural-language approaches separately, emphasizing how their reasoning methods and evaluation protocols differ.
@@ 4.1.1 Natural Language Hypothesis Generation with LLMs
Prompt-Based Methods: Due to the lack of large-scale, domain-specific data for hypothesis generation, most abduction approaches rely on prompt-based methods that are easy to deploy and don’t require extensive additional data. For instance, when provided with observations expressed in natural language and asked to generate a plausible hypothesis that explains them, both Wiegreffe et al. (2022) and Qi et al. (2024) employ few-shot prompting to guide LLMs in generating hypotheses. Specifically, Wiegreffe et al. (2022) constructs few-shot examples using a triplet format (question, answer, explanation). In solving a task of generating biomedical hypotheses with given observations, Qi et al. (2024) embeds a small set of independent observation-to-hypothesis pairs in the prompt. By showing how each block of biomedical background observations maps to its corresponding hypothesis, the model learns to extract relevant domain cues and generate novel biomedical hypotheses. Their findings indicate that including more examples in the prompt tends to reduce the novelty of the generated hypotheses while increasing their correctness. Furthermore, Yang et al. (2024a) propose a pipeline for hypothesis generation that involves five prompt-based modules: one to generate hypotheses, one to test deductive consistency, one to verify that the hypothesis is not merely a copy of the given context, one to assess its generalizability, and one to determine whether the hypothesis is trivial.
@@ 5.2 Evaluation for Hypothesis Application
Building on this idea, Yu et al. (2024b) create the TURTLEBENCH benchmark, inspired by the “Turtle Soup” game in which players deduce a story’s hidden explanation by asking yes/no questions; in TURTLEBENCH, the LLM instead answers human-annotated questions with “True,” “False,” or “Not Relevant,” across 1,532 high-quality question-answer pairs sourced from an online platform to test whether it can fully follow a story and provide accurate answers. Similarly, Mu et al. (2024) introduced the RULES benchmark, comprising 14 rule-following scenarios, each paired with concise test cases and programmatic evaluation functions that objectively assess adherence to specified rules. In addition, Cai et al. (2025) presented the Holiday Puzzle benchmark, which features multiple holiday schedule scenarios ranging from simple single-week planning to multi-phase arrangements and complex date arithmetic tasks, again using test cases and evaluation functions to verify correct computation of extra holiday rest days under provided rules. Moreover, Sun et al. (2024) constructed RuleBench to evaluate not only whether models can produce correct answers based on factual rules but also their ability to apply counterfactual rules, designed to yield incorrect outcomes, and experiments show that while LLMs achieve near-perfect accuracy on factual rules, their performance drops dramatically under counterfactual rules, revealing a significant gap in counterfactual rule-following capability.
@@ 6 Hypothesis Validation
According to Peirce, induction validates a hypothesis by updating its confidence when new evidence appears. However, in studies that focus exclusively on induction, tasks are typically one-off: a hypothesis (or set of hypotheses) and a collection of observations are provided, and there is no iterative updating of confidence. A simplified framework for hypothesis validation treats it as a multiple-choice problem: given observations $\mathbb{O}=\{o_{1},o_{2},\dots,o_{n}\}$ and a set of hypothesis =$\mathbb{H}=\{h_{1},h_{2},\dots,h_{m}\}$, the model selects the most possible hypothesis. In simpler scenarios, where only one hypothesis is provided, the model determines whether the hypothesis correctly explains the observations. In the next section, when combined with deduction and abduction, induction can subsequently be used to iteratively update the confidence in the hypothesis.
Natural language representations add significant complexity to induction. In formal language settings, all necessary information is explicitly provided, and reasoning follows rigorous, well-defined steps. In contrast, validating a natural language hypothesis often requires commonsense knowledge and interpretation of nuanced language. For example, consider the observations “Neil wanted to see the mountains of Asia” and “Neil loved being so close to the mountains in Nepal,” with candidate hypotheses “Neil booked a trip online” and “Neil took a trip to see the Rocky Mountains instead.” Here, the nuanced meaning of the term “instead” and the geographic relationships require careful analysis and may lead to different conclusions. Indeed, Zhang et al. (2020) reports that, when verifying their dataset where five annotators judged the plausibility of handwritten hypotheses, disagreements occurred in 62.34% of 1,365 explanations, underscoring the challenge of natural language hypothesis validation.
@@ 6.2.2 Natural language Evaluation
Bhagavatula et al. (2020) introduce the ART benchmark, comprising roughly 20k narrative contexts where each sample includes two time-ordered observations, one depicting a story’s start ($o_{1}$) and the other its outcome ($o_{2}$), alongside two hypotheses: a plausible explanation ($h^{+}$) and a less plausible one ($h^{-}$), challenging models to choose the best explanatory hypothesis and enabling adaptation to hypothesis-generation tasks evaluated against ground-truth explanations. Similarly, Jiang et al. (2023) present the BRAINTEASER benchmark of about 1.1k lateral-thinking puzzles, each offering a question with multiple-choice answers, one that defies commonsense and several conventional distractors, in both sentence (narrative) and word (meaning-alteration) formats to test creative reasoning, with additional semantic and context reconstruction variants assessing reasoning consistency and robustness across formulations. Moreover, Del & Fishel (2023) introduced the True Detective benchmark for deep hypothesis validation, featuring 191 long-form detective puzzles ($\approx$ 1200 words each) from the “5 Minute Mystery” platform, where models (and humans) select the correct explanation from 4–5 options, human accuracy averages 47%, top solvers exceed 80%, and each puzzle includes golden chain-of-thought explanations detailing the reasoning steps that lead to the correct answer.
@@ 7.1 Passive Hypothesis Discovery
Both HypoGeniC and HtT simplify hypothesis discovery by relying on benchmark questions that include ground-truth answers. This configuration allows an external algorithm, not the LLMs themselves, to validate generated hypotheses and update their confidence based on the correctness of predictions. In real-world scenarios, where no ground-truth answers are available, these frameworks become inapplicable and would require substantial adaptation.

### v2

@@ 1 Introduction
[…]
Reasoning for Hypothesis Discovery
Hypothesis Generation w/ Abduction § 4 Method § 4.1 Natural Language Hypothesis Prompt Method Wiegreffe et al. (2022); Qi et al. (2024); Yang et al. (2024a) RAG Method Hu et al. (2024); Yang et al. (2025; 2024b); Xiong et al. (2024) Chai et al. (2024) Human supported Zhao et al. (2024); Si et al. (2024); Pu et al. (2025) Formal Language Hypothesis W/ Formal observation Young et al. (2022); Nguyen et al. (2023) W/ NL observation Cheng et al. (2024); Wang et al. (2024); Qiu et al. (2024) il Lee et al. (2025); Movva et al. (2025) Evaluation § 4.2 Natural Language Hypothesis Prediction Based Sinha et al. (2019); Weston et al. (2015); Balepur et al. (2024) Shi et al. (2023); Wang et al. (2025) Ground Truth Based Yang et al. (2024a); Qi et al. (2024); Movva et al. (2025) Yang et al. (2024b) Human Evaluation Zhao et al. (2024); Yang et al. (2024b); Hu et al. (2024); Qi et al. (2024) Yang et al. (2025) Formal Language Hypothesis Prediction Based Rule (2020); Chollet (2019); Liu et al. (2024); Li et al. (2025a) Chen et al. (2024) Ground Truth Based Bowen et al. (2024); Hua et al. (2025); Young et al. (2022) Hypothesis Application w/ Deduction § 5 Method § 5.1 Natural Language Hypothesis LLM as FL Parser Pan et al. (2023); Olausson et al. (2023); Kalyanpur et al. (2024) Fine-Tune Based Sun et al. (2024) Prompt-Based Sun et al. (2024); Cai et al. (2025); Ling et al. (2023) Formal Language Hypothesis N/A Evaluation § 5.2 Natural Language Hypothesis Prediction Based Yu et al. (2024b); Mu et al. (2024); Cai et al. (2025); Sun et al. (2024) Formal Language Hypothesis N/A Hypothesis Validation w/ Induction § 6 Method § 6.1 Natural Language Hypothesis Prompt-Based Lampinen et al. (2022); Sun et al. (2024) Fine-Tune Based Zhao et al. (2023); Chan et al. (2023); Sun et al. (2024) Formal Language Hypothesis Prompt-Based He & Lu (2024) Evaluation § 6.2 Natural Language Hypothesis True/False Based Lampinen et al. (2022); Rudinger et al. (2020); Zhang et al. (2025) Selection Based Bhagavatula et al. (2020); Jiang et al. (2023); Del & Fishel (2023) Formal Language Hypothesis Ground Truth Based He & Lu (2024) Entire Hypothesis Discovery § 7 Passive Hypothesis Discovery § 7.1 Zhu et al. (2024); Zhou et al. (2024) Proactive Hypothesis Discovery § 7.2 Li et al. (2024); Jung et al. (2022); Seals & Shalin (2024) Conti et al. (2024) Hypothesis Discovery with Real World Simulation § 7.3 Xu et al. (2023); Wang et al. (2022a); Jansen et al. (2024) He et al. (2024)
@@ 2.2 Reasoning
Deductive Reasoning is the process of logically deriving specific conclusions from general hypotheses or rules. If the initial hypotheses are true, deduction guarantees that the derived conclusions must also be true. For instance, from the general rule “All swans are white” and the observation “This bird is a swan,” we logically conclude “This bird must be white”. While traditional deductive reasoning tasks, such as instruction-following and standard problem-solving, have been extensively studied with LLMs (Pan et al., 2023; Wei et al., 2022; Liu et al., 2025; Huang & Chang, 2023), deductive reasoning in the context of hypothesis discovery poses unique challenges. Specifically, it emphasizes inferential rule-following, requiring models to consistently apply hypotheses or rules to derive new and potentially unfamiliar conclusions, even when these hypotheses are counterfactual, unfamiliar, or incorrect. For example, when a flawed hypothesis is introduced in an unfamiliar domain, inferential rule‐following requires us to strictly derive its predicted consequence, even if that consequence itself is incorrect. By comparing this consequence with experimental data, we can directly assess the hypothesis’s validity and guide its revision. Conversely, if the deductive process is unreliable, we may overlook real contradictions and thus retain invalid hypotheses or discard valid ones. Indeed, recent work shows that although LLMs can demonstrate strong deductive performance on in-distribution tasks, they rely heavily on surface‐level pattern matching and fail to generalize their inferential rule‐following to novel or counterfactual scenarios (Pu et al., 2025; Mirzadeh et al., 2024; Kang et al., 2024; Yan et al., 2025).
@@ 2.3 Rule Representation: Formal Language vs Natural Language
Natural Language
English: ‘Sam is a dragon”
English commonsense
of ‘dragon”
Sam is dangerous
Chinese: ‘Sam is a dragon”
Chinese commonsense
of ‘dragon”
Sam brings good fortune
and a bountiful harvest
Formal Language
FOL: $\mathrm{Dragon}(\mathrm{Sam})$
$\forall x(\mathrm{Dragon}(x)\to\mathrm{Fly}(x))$
$\mathrm{Fly(Sam)}$
Code: Sam = Dragon()
Class Dragon:
   def fly(self):
Sam.fly()
There are many ways to represent hypotheses and rules, which we broadly divide into two categories: formal languages (FL) and natural languages (NL). Formal languages, such as first‑order logic and programming languages, are systematic and rule‑bound. After real‑world entities are encoded as explicit literals, precise inference rules yield provably correct and sound conclusions, making these systems well suited to deductive reasoning. Yet the encoding process strips away many subtle semantic relationships and commonsense knowledge, limiting the system’s ability to handle the creative, defeasible reasoning required for abduction and induction (McCarthy & Hayes, 1981; Reiter, 1980; Hanks & McDermott, 1987; Liu et al., 2025; Yu et al., 2024a; Huang & Chang, 2023). Natural language preserves those nuances and aligns more closely with human cognition, so it is better suited to abductive and inductive tasks. However, its meanings are implicit and context‑dependent, making it difficult to define a deterministic reasoning pipeline and reducing the reliability of the resulting inferences (See in Table  1). Accordingly, the following sections treat formal‑language and natural‑language approaches separately, emphasizing how their reasoning methods and evaluation protocols differ.
@@ 4.1.1 Natural Language Hypothesis Generation with LLMs
Prompt-Based Methods: Due to the lack of large‑scale, domain‑specific data for hypothesis generation, most abduction approaches rely on prompt‑based methods that are easy to deploy and don’t require extensive additional data. For instance, when provided with observations expressed in natural language and asked to generate a plausible hypothesis that explains them, both Wiegreffe et al. (2022) and Qi et al. (2024) employ few-shot prompting to guide LLMs in generating hypotheses. Specifically, Wiegreffe et al. (2022) constructs few-shot examples using a triplet format (question, answer, explanation). In solving a task of generating biomedical hypotheses with given observations, Qi et al. (2024) embeds a small set of independent observation‑to‑hypothesis pairs in the prompt. By showing how each block of biomedical background observations maps to its corresponding hypothesis, the model learns to extract relevant domain cues and generate novel biomedical hypotheses. Their findings indicate that including more examples in the prompt tends to reduce the novelty of the generated hypotheses while increasing their correctness. Furthermore, Yang et al. (2024a) propose a pipeline for hypothesis generation that involves five prompt-based modules: one to generate hypotheses, one to test deductive consistency, one to verify that the hypothesis is not merely a copy of the given context, one to assess its generalizability, and one to determine whether the hypothesis is trivial.
@@ 5.2 Evaluation for Hypothesis Application
Building on this idea, Yu et al. (2024b) create the TURTLEBENCH benchmark, inspired by the “Turtle Soup” game in which players deduce a story’s hidden explanation by asking yes/no questions; in TURTLEBENCH, the LLM instead answers human‑annotated questions with “True,” “False,” or “Not Relevant,” across 1,532 high‑quality question‑answer pairs sourced from an online platform to test whether it can fully follow a story and provide accurate answers. Similarly, Mu et al. (2024) introduced the RULES benchmark, comprising 14 rule‑following scenarios, each paired with concise test cases and programmatic evaluation functions that objectively assess adherence to specified rules. In addition, Cai et al. (2025) presented the Holiday Puzzle benchmark, which features multiple holiday schedule scenarios ranging from simple single‑week planning to multi‑phase arrangements and complex date arithmetic tasks, again using test cases and evaluation functions to verify correct computation of extra holiday rest days under provided rules. Moreover, Sun et al. (2024) constructed RuleBench to evaluate not only whether models can produce correct answers based on factual rules but also their ability to apply counterfactual rules, designed to yield incorrect outcomes, and experiments show that while LLMs achieve near‑perfect accuracy on factual rules, their performance drops dramatically under counterfactual rules, revealing a significant gap in counterfactual rule‑following capability.
@@ 6 Hypothesis Validation
According to Peirce, induction validates a hypothesis by updating its confidence when new evidence appears. However, in studies that focus exclusively on induction, tasks are typically one‑off: a hypothesis (or set of hypotheses) and a collection of observations are provided, and there is no iterative updating of confidence. A simplified framework for hypothesis validation treats it as a multiple-choice problem: given observations $\mathbb{O}=\{o_{1},o_{2},\dots,o_{n}\}$ and a set of hypothesis $\mathbb{H}=\{h_{1},h_{2},\dots,h_{m}\}$, the model selects the most possible hypothesis. In simpler scenarios, where only one hypothesis is provided, the model determines whether the hypothesis correctly explains the observations. In the next section, when combined with deduction and abduction, induction can subsequently be used to iteratively update the confidence in the hypothesis.
Natural language representations add significant complexity to induction. In formal language settings, all necessary information is explicitly provided, and reasoning follows rigorous, well-defined steps. In contrast, validating a natural language hypothesis often requires commonsense knowledge and interpretation of nuanced language. For example, consider the observations “Neil wanted to see the mountains of Asia” and “Neil loved being so close to the mountains in Nepal,” with candidate hypotheses “Neil booked a trip online” and “Neil took a trip to see the Rocky Mountains instead.” Here, the nuanced meaning of the term “instead” and the geographic relationships require careful analysis and may lead to different conclusions. Indeed, Zhao et al. (2023) reports that, when verifying their dataset where five annotators judged the plausibility of handwritten hypotheses, disagreements occurred in 62.34% of 1,365 explanations, underscoring the challenge of natural language hypothesis validation.
@@ 6.2.2 Natural language Evaluation
Bhagavatula et al. (2020) introduce the ART benchmark, comprising roughly 20k narrative contexts where each sample includes two time‑ordered observations, one depicting a story’s start ($o_{1}$) and the other its outcome ($o_{2}$), alongside two hypotheses: a plausible explanation ($h^{+}$) and a less plausible one ($h^{-}$), challenging models to choose the best explanatory hypothesis and enabling adaptation to hypothesis‑generation tasks evaluated against ground‑truth explanations. Similarly, Jiang et al. (2023) present the BRAINTEASER benchmark of about 1.1k lateral‑thinking puzzles, each offering a question with multiple‑choice answers, one that defies commonsense and several conventional distractors, in both sentence (narrative) and word (meaning‑alteration) formats to test creative reasoning, with additional semantic and context reconstruction variants assessing reasoning consistency and robustness across formulations. Moreover, Del & Fishel (2023) introduced the True Detective benchmark for deep hypothesis validation, featuring 191 long‑form detective puzzles ($\approx$ 1200 words each) from the “5 Minute Mystery” platform, where models (and humans) select the correct explanation from 4–5 options, human accuracy averages 47%, top solvers exceed 80%, and each puzzle includes golden chain‑of‑thought explanations detailing the reasoning steps that lead to the correct answer.
@@ 7.1 Passive Hypothesis Discovery
Both HypoGeniC and HtT simplify hypothesis discovery by relying on benchmark questions that include ground‑truth answers. This configuration allows an external algorithm, not the LLMs themselves, to validate generated hypotheses and update their confidence based on the correctness of predictions. In real‑world scenarios, where no ground‑truth answers are available, these frameworks become inapplicable and would require substantial adaptation.